From GlueXWiki
Jump to: navigation, search

March 19, 2015 FDC+CDC meeting


  1. Instructions for Bluejeans meeting connection
  2. FDC meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. Gas issues
  2. CDC update (Beni, Mike)[5]
  3. FDC update
  4. Electronics: logbook entry 3325323, 3325070 (Fernando, Cody, Chris, Nick)
  5. Engineering (Dave)
  6. Other


Participants: Naomi (CMU), Sean (NU), Fernando, Dave, Chris, Nick, Beni, Sergey, and Lubomir (JLab).

Gas issues

- Lubomir presented summary of the FDC gas gain measurements with 55Fe source: first plot shows charge distributions at different HV in the Hall, package 4. Second plot is a summary of all the measurements at 2200/-500V both in the Hall and in the EEL126 with the spare package, as a function of the atmospheric pressure. 1% change of the pressure results in 5-7% change of the gain, as expected from the literature. By some reasons the charge on the upper strips is bigger than on the down ones for both measurements in the EEL and in the Hall; needs further studies. Third plot shows the gain-HV dependence at fixed (+/- 0.2 kPa) atmospheric pressure from the measurements and also from Garfield simulations. Garfield predicts somewhat stronger dependence: Sergey explains it with space charge effects at higher gains when using 55Fe source. Assuming a mixture of 45/55 Ar/CO2 (plus 0.3% alcohol) in the Hall, Garfield predicts higher gains, all this relative to the standard gas mixture 40/60 in the EEL. A simple explanation might be a difference of the gas temperature in the Hall and in the EEL: 6 degC lower temperature in the Hall will explain it, or 1% (absolute) change of the CO2. Conclusion: generally the gain measurements in the EEL and in the Hall are in agreement with a change of the CO2 from 60% to 55% (+/-1%). Last plot shows previous gain measurements for four gas mixtures using different techniques.

- Beni continued changing different configurations in the gas mixing system trying to identify the source of the disagreement without success so far: the biggest change in the max drift time was 8% but was not repeatable.

CDC update

- Mike showed a plot with the positions of the noisy channels: most are at ~7 o'clock. The HV filter was installed on the channels at ~9 o'clock and this sector is indeed very clean.

FDC update

- The chiller is consuming a lot of Fluorinert. After 10 days of operation we had to refill it with ~1.5 galons (see photos attached). Need to sniff the connections and later possibly the chiller inside. As discussed with Eugene and Tim, the plan is to buy another chiller like this and connected it to the FDC right away.

- The 8 wire channels with no signals reported on the last meeting were fixed: the connector was not properly connected at the front of the TDC.

- Picture: no corrosion at all on one of the cathodes on the spare package.


- HV filters will come by the end of the month, so hopefully we will have a good data in April. Nick needs and will take from the Hall the only filter that is now installed on the CDC HV.


- Dave, Lubomir discussed with Bill possible replacement of the CO2 sensor pump. Bill contacted a company with long lasting pumps however they all work on flows > several liters per minute. We decided to go with a diaphragm pump that has ~1 year life with a variable flow (spec attached): we need 0.1 l/min for FDC and 0.1-0.5 l/min for the CDC. Fernando agreed to install one MPOD in a BCAL crate to vary the voltage of the pump.